OpenStudy (anonymous):

HELP is x^3 a function, but not a one to one function?

jimthompson5910 (jim_thompson5910):

f(x) = x^3 is a function that is also one-to-one

jimthompson5910 (jim_thompson5910):

if f(a) = f(b) leads to a = b, then f(x) is a one-to-one function, so let's prove that

jimthompson5910 (jim_thompson5910):

f(x) = x^3 f(a) = a^3 ----------- f(x) = x^3 f(b) = b^3 ----------- f(a) = f(b) a^3 = b^3 a = b ... take the cube root of both sides So this proves that f(x) = x^3 is a one-to-one function

OpenStudy (anonymous):

Ah, thank you!

jimthompson5910 (jim_thompson5910):

np